evaluate the following:


1. (i^41+1/i^257)^9 ?

Grade:10

1 Answers

Akash Verma
11 years ago

i^41=i^2
1/i^257=i/i^258=i
so,
the question simplifies to:
(i^2+i)^9=(-1)^9=-1
so the answer is -1